∫ (sec θ - tan θ)² dθ


1
Expert's answer
2021-04-15T17:21:55-0400

solve "\u222b (sec \u03b8 - tan \u03b8)\u00b2 d\u03b8"

"=\u222b(sec^2-2 sec\u03b8tan\u03b8+tan^2 \u03b8)d\u03b8"

"=\u222b(1+tan^2 \u03b8-(2sec \u03b8tan\u03b8+tan^2 \u03b8)d\u03b8"

"=\u222b(1-2sec\u03b8tan\u03b8+2tan^2 \u03b8)d\u03b8"

"=\u222bd\u03b8-2\u222b(2sec\u03b8tan\u03b8d\u03b8)+2\u222b(tan^2 \u03b8d\u03b8)"

"=\u03b8-2\u222bsec\u03b8tan\u03b8d\u03b8+\u222btan^2 \u03b8d\u03b8"

for

"\u222bsec\u03b8tan\u03b8d\u03b8, let u=sec\u03b8 , du=sec\u03b8tan\u03b8d\u03b8"

"\u061e \u222bsec\u03b8tan\u03b8d\u03b8=\u222bdu=u\n =sec\u03b8\t\u2026(ii)"

and for

"\int tan ^2 \theta d\theta" we define "ǀn =∫tan^2 θdθ=\frac {tan^{-1} θ}{n-1}-ǀ_{n-2}"

as the reduction formula

Thus

"ǀ_2=∫tan^2 θdθ= \frac {tanθ}{1} -ǀ_0"

"=tan\u03b8-\u03b8\u2026(iii)"

Replacing (ii) and (iii) in (i)gives,

"=\u222b(sec\u03b8-tan\u03b8)^2 d\u03b8"

"=\u03b8-2 sec \u03b8+2(tan\u03b8-\u03b8)+c"

"=2tan\u03b8-2sec\u03b8-\u03b8+c"